Pappy's Toys makes two models of a metal toy—Standard and DeLuxe. Both models are produced on a single machine. The price and costs of the two models appear as follows.

Standard DeLuxe
Price per unit $ 100 $ 120
Variable cost per unit $ 44 $ 48
Machine hours per unit 0.5 1.5

The one machine that used to produce both models has a capacity of 102,000 machine-hours per quarter. Fixed manufacturing costs per quarter are $860,000.

Required:

a. Suppose that the maximum unit sales in a quarter that Pappy's can achieve is 224,000 units of each product. How many units of each model should Pappy's produce in a quarter?

b. Suppose that the maximum unit sales in a quarter that Pappy's can achieve is 126,000 units of each product. How many units of each model should Pappy's produce in a quarter?
Answer #1
Standard Deluxe
Price per unit 100 120
Less: Variable cost per unit 44 48
Contribution margin per unit 56 72
Divide by Machine hours per unit 0.5 1.5
Contribution margin per machine hour 112 48
Standard model should be produced first
a
Units of Standard model 204000 =102000/0.5
Standard model 204000 units
Deluxe model 0 units
b
Machine hours for Standard model 63000 =126000*0.5
Machine hours remaining for Deluxe model 39000 =102000-63000
Units of Deluxe model 26000 =39000/1.5
Standard model 126000 units
Deluxe model 26000 units